How the numbers are being divided by avagadro number show the division
pls show it

Grade:9

1 Answers

swarup kumar majee
6 years ago
avogardo number=6.022*1023
1 mole of any particle=6.022*1023 number of particles
for example 1 mole of oxigen atom=6.022*1023atoms